# Corresponding Author

### v3

@@ 5.1 Personalized Object Generation
An important branch in TTF methods is to optimize the learnable token embeddings. The first work started from Textual Inversion 1gal2022image , which applies a simple yet effective method that introduces the unique modifier as a new token to represent the SoI. One of the significant benefits of this method is its minimal storage requirement as the new tokens consumes just a few kilobytes. However, the method compresses complex visual features into a small set of parameters, which can lead to long convergence times and a potential loss in visual fidelity. Recent work aims to address these limitations through several directions. DVAR 69voronov2024loss improves training efficiency by proposing a clear stopping criterion by removing all randomness to indicate the convergence. For enhanced representation, P+ 23voynov2023p+ introduces distinct learnable tokens across different layers of the U-Net architecture, thereby offering better attribute control through additional learnable parameters. NeTI 61alaluf2023neural advances this concept by proposing a neural mapper that adaptively outputs token embeddings based on the denoising timestep and specific U-Net layers. ProSpect 30zhang2023prospect recommends optimizing multiple token embeddings tailored to different denoising timesteps based on the observation that different types of prompts, like layout, color, structure, and texture, are activated at different stages of the denoising process. Similarly, a study by 79agarwal2023image shows layered activation insight to learn distinct attributes by selectively activating the tokens within their respective scopes. Later, HiFiTuner 48wang2023hifi integrates multiple techniques into the learnable token, including mask-guided loss function, parameter regularization, time-dependent embedding, and generation refinement assisted by the nearest reference. Alternative approaches, like DreamArtist 68dong2022dreamartist , choose to optimize both negative and positive prompt embeddings to refine the detail preservation. In addition to these token-level refinement approaches, the field continues to evolve with novel techniques such as InstructBooth’s reinforcement learning framework 56chae2023instructbooth and gradient-free evolutionary optimization 73fei2023gradient . In summary, recent developments following the foundational work of Textual Inversion focus on reducing training times and enhancing the visual quality of generated images.
@@
In the realm of TTF methods for PCS, there is a clear shift towards fine-tuning model weights rather than just the token embeddings. This approach often addresses the limitations where token embeddings alone struggle to capture complex semantics uncovered in the pre-training data 46gu2024mix ; 80zhao2023catversion . DreamBooth 2ruiz2023dreambooth proposes to use a unique modifier by a rare token to represent the SoI and fine-tune the whole parameters of the diffusion model. Besides, a regularization dataset containing 20-30 images with the same category as SoI is adopted to overcome the overfitting problem. These two combined approaches achieve impressive performance that largely promotes the progress of the research on image personalization. However, fine-tuning the entire model for each new object causes considerable storage costs, potentially rendering widespread application. To address this, Custom Diffusion 6kumari2023multi focuses on identifying and fine-tuning critical parameters, particularly the key-value projections in cross-attention layers, to achieve a balance of visual fidelity and storage efficiency. Further approach, Perfusion 7tewel2023key , also adopts the cross-attention fine-tuning and proposes to regularize the update direction of the K (key) projection towards the super-category token embedding and the V (value) projection towards the learnable token embedding. COMCAT 75xiao2023comcat introduces a low-rank approximation of attention matrices, which drastically reduces storage requirements to 6 MB while maintaining high fidelity in the outputs. Additionally, methods like adapters 72xiang2023closer ; 5sohn2024styledrop and LoRA variants 46gu2024mix ; 58ruiz2023hyperdreambooth ; 115kong2024omg ; 51lu2024object ; 95achlioptas2023stellar ; 76yeh2023navigating are increasingly utilized in personalized generation for parameter-efficient fine-tuning. It is worth noting that the token embedding fine-tuning is compatible with diffusion weight fine-tuning. Multiple methods 90hyung2023magicapture ; 74avrahami2023break ; 86hua2023dreamtuner have started using combined weight fine-tuning.
@@
PTA framework. For practical deployment of PCS systems, fast response time is a crucial factor. PTA methods enable real-time generation (less than 10 sec/subject) by leveraging large-scale pre-training to avoid per-subject optimization during the inference phase. Re-Imagen 67chen2022re introduces a retrieval-augmented generative approach, which leverages features from text-image pairs retrieved via a specific prompt. While it is not specifically tailored for object personalization, it demonstrates the feasibility of training reference-conditioned frameworks. Later, ELITE 19wei2023elite specifically targets image personalization by combining the global reference features with text embedding while incorporating local features that exclude irrelevant backgrounds. Both fused features and local features serve as conditions for the denoising process. Similarly, InstantBooth 22shi2023instantbooth retrains CLIP models to extract image features and patch features, which are injected into the diffusion model via the attention mechanism and learnable adapter, respectively. Additionally, UMM-Diffusion 71ma2023unified designs a multi-modal encoder that produces fused features based on the reference image and text prompt. The text features and multi-modal hidden state are seen as guidance signals to predict a mixed noise. Another work, SuTI 11chen2024subject , adopts the same architecture as Re-Imagen. The difference lies in the training samples which are produced by a massive number of TTF models, each tuned on a particular subject set. This strategy promotes a more precise alignment with personalization at an instance level rather than the class level of Re-Imagen. Moreover, Domain-Agnostic 17arar2023domain combines a contrastive-based regularization technique to push the pseudo embedding produced by the image encoder towards the existing nearest pre-trained token. Besides, They introduce a dual-path attention module separately conditioned on the nearest token and pseudo embedding. Compared to the methods that use separate encoders to process a single modality, some works have explored the usage of pre-trained multi-modal large language models (MLLM) that can process text and image modality within a unified framework. For example, BLIP-Diffusion 10li2024blip utilizes the pre-trained BLIP2 li2023blip that encodes multimodal inputs including the SoI reference and a class noun. The output embedding is then concatenated with context description and serves as a condition to generate images. Further, Customization Assistant 84zhou2023customization and KOSMOS-G 77pan2023kosmos replace the text encoder of Stable Diffusion with a pre-trained MLLM to output a fused feature based on the reference and context description. Meanwhile, to meet the standard format of Stable Diffusion, a network is trained to align the dimension of the output embedding.
@@ 5.2 Personalized Style Generation
TTF framework. In this field, StyleDrop 5sohn2024styledrop leverages adapter tuning to efficiently capture the style from a single reference image. This method demonstrates the effectiveness through iterative training, utilizing synthesized images refined by feedback mechanisms, like human evaluations and CLIP scores. This approach not only enhances style learning but also ensures that the generated styles align closely with human aesthetic judgments. Later, GAL 151zhang2024generative proposes an uncertainty-based evaluation strategy to filter high-quality synthetic-style data and uses a weighted schema to balance the contribution of the additional samples and the original reference. Furthermore, StyleAligned 98hertz2023style focuses on maintaining stylistic consistency across a batch of images. This is achieved by using the first image as a reference, which acts as an additional key and value in the self-attention layers, ensuring that all subsequent images in the batch adhere to the same stylistic guidelines. Style-friendly 211choi2024style introduces a novel diffusion model fine-tuning approach that enhances personalized artistic style generation by adaptively biasing noise sampling toward higher noise levels, where stylistic features emerge.
PTA framework. For the PTA framework, StyleAdapter 26wang2023styleadapter employs a dual-path cross-attention mechanism within the PTA framework. This model introduces a specialized embedding module designed to extract and integrate global features from multiple style references. Diptych Prompting 209shin2024large utilizes an inpainting mechanism to draw another image with the same style of the reference part.
@@ 5.3 Personalized Face Generation
Personalized face generation aims to generate diverse ID images that adhere to text prompt specifications, utilizing only a few initial face images. Compared to general object personalization, the scope is narrowed to a specific class, humans. It is easy to acquire large-scale human-centric datasets zheng2022general ; yang2020facescape ; cao2018vggface2 and utilize pre-trained models in well-developed areas, like face landmark detection wu2019facial ; khabarlak2021fast and face recognition adjabi2020past .
TTF framework. As for the TTF methods, PromptNet 55zhou2023enhancing trains a diffusion-based network that encodes the input image and noisy latent to a word embedding. To alleviate the overfitting problem, the noises predicted by the word embedding and context description are balanced through fusion sampling through classifier-free guidance. HyperDreamBooth 58ruiz2023hyperdreambooth proposes a second-time fine-tuning strategy to further enhance the fidelity after training a PTA model on a large-scale dataset. Additionally, Celeb Basis 60yuan2023inserting provides a novel idea that the personalized ID can be viewed as the composition of celebrity’s face which has been learned by the pre-trained diffusion model. Based on this hypothesis, a simple MLP is optimized to transform face features into the celeb embedding space.
PTA framework. Thanks to the abundance of available datasets featuring the same individual in various contexts, which provide valuable data for pre-training PTA methods, the number of works in PTA frameworks is rapidly increasing. Face0 88valevski2023face0 crops the face region to extract refined embeddings and concatenates them with text features. During the sampling phase, the output of classifier-free guidance is replaced by a weighted combination of the noise patterns predicted by face-only embedding, text-only embedding, and concatenated face-text embedding. The $\mathcal{W}+$ Adapter 91li2023stylegan constructs a mapping network and residual cross-attention modules to transform the facial features from the StyleGAN karras2019style $\mathcal{W}+$ space into the text embedding space of Stable Diffusion. FaceStudio 31yan2023facestudio adapts the cross-attention layer to support hybrid guidance including stylized images, facial images, and textual prompts. Moreover, PhotoMaker 28li2023photomaker constructs a high-quality dataset through a meticulous data collection and filtering pipeline. They use a two-layer MLP to fuse ID features and class embeddings for an overall representation of human portrait. PortraitBooth 96peng2023portraitbooth also employs a simple MLP, which fuses the text condition and shallow features of a pre-trained face recognition model. To ensure expression manipulation and facial fidelity, they add another expression token and incorporate the identity preservation loss and mask-based cross-attention loss. InstantID 37wang2024instantid additionally introduces a variant of ControlNet that takes facial landmarks as input, providing stronger guiding signals compared to the methods that solely rely on attention fusion.
@@ 5.4 Multiple Subject Composition
TTF framework. This task presents a challenge for the TTF methods, particularly in how to integrate the parameters within the same module which are separately fine-tuned for individual SoI. Some works focus on the one-for-one generation following a fusion mechanism. For instance, Custom Diffusion 6kumari2023multi proposes a constrained optimization method to merge the cross-attention key-value projection weights with the goal of maximizing reconstruction performance for each subject. Mix-of-Show 46gu2024mix fuses the LoRA hu2021lora weights with the same optimization objective. StyleDrop 5sohn2024styledrop dynamically summarizes noise predictions from each personalized diffusion model. In OMG 115kong2024omg , the latent predicted by each LoRA-tuned model is spatially composited using the subject mask. Joint training is another strategy to cover all expected subjects. SVDiff 8han2023svdiff employs a data augmentation method called Cut-Mix to compose several subjects together and applies a location loss to regularize attention maps, ensuring alignment between each subject and its corresponding token. Similar strategies are found in other works 6kumari2023multi ; 60yuan2023inserting which train a single model by reconstructing the appearance of every SoI. There are advanced control mechanisms designed to manage multiple subjects. Cones 4liu2023cones proposes to find a small cluster of neurons that preserve the most information about SoI. The neurons belonging to different SoI will be simultaneously activated to generate the combination. Compositional Inversion 9zhang2023compositional introduces spatial region assignment to different subjects to improve the composition success rate.
PTA framework. For PTA frameworks, multi-subject generation is achieved through specialized architectural designs. Fastcomposer 54xiao2023fastcomposer , Subject-Diffusion 16ma2023subject , and $\lambda$-eclipse 222patel2024lambda place each subject feature in its corresponding placeholder in the text embedding, ensuring a seamless and efficient combination. CustomNet 29yuan2023customnet and MIGC 174zhou2024migc train a PTA network that supports location control for each subject. SSR-Encoder 221zhang2024ssr implements an encoder to selectively preserve the desired subject feature and a cross-attention module to support multi-subject feature fusion.
@@ 5.5 High-level Semantic Personalization
TTF framework. For now, all researches in this field are based on the TTF framework. ReVersion 63huang2023reversion intends to invert object relations from references. Specifically, they use a contrastive loss to guide the optimization of the token embedding towards specific clusters of Part-of-Speech tags, such as prepositions, nouns, and verbs. Meanwhile, they also increase the likelihood of adding noise at larger timesteps during the training process to emphasize the extraction of high-level semantic features. Lego 45motamed2023lego focuses on more general concepts, such as adjectives, which are frequently intertwined with the subject appearance. The concept can be learned from a contrastive loss applied to the dataset comprising clean subject images and images that embody the desired adjectives. Moreover, ADI 43huang2023learning aims to learn the action-specific identifier from the references. To ensure the inversion only focuses on the desired action, ADI extracts gradient invariance from a constructed triplet sample and applies a threshold to mask out the irrelevant feature channels.
@@ 6.1 Personalization on Extra Conditions
Recent personalization tasks tend to include additional conditions for diverse content customization. One common application is to condition on a fixed source image. For example, PhotoSwap 15gu2024photoswap introduces a new task that replaces the subject in the source image with the SoI from the reference image. To address this requirement, they first fine-tunes a diffusion model on the references to obtain a personalized model. To preserve the background of the source image, they initialize the noise with DDIM inversion song2020denoising and replace the intermediate feature maps with those derived from source image generation during inference time. Later, MagiCapture 91li2023stylegan broadens the scope to face customization. Another similar application can be found in Virtual Try-on, which aims to fit selected clothing onto a target person. The complexities of this task have been thoroughly analyzed in another survey islam2024deep .
Additional conditions in personalization tasks may include adjusting the layout 152chen2024training , transforming sketches 94chen2023democaricature , controlling viewpoint 29yuan2023customnet ; 163kumari2024customizing , or modifying poses 37wang2024instantid . Each of these conditions presents unique challenges and requires specialized approaches to integrate these elements seamlessly into the personalized content.

### v4

@@ 5.1 Personalized Object Generation
An important branch in TTF methods is to optimize the learnable token embeddings. The first work started from Textual Inversion [7], which applies a simple yet effective method that introduces the unique modifier as a new token to represent the SoI. One of the significant benefits of this method is its minimal storage requirement as the new tokens consumes just a few kilobytes. However, the method compresses complex visual features into a small set of parameters, which can lead to long convergence times and a potential loss in visual fidelity. Recent work aims to address these limitations through several directions. DVAR [90] improves training efficiency by proposing a clear stopping criterion by removing all randomness to indicate the convergence. For enhanced representation, P+ [18] introduces distinct learnable tokens across different layers of the U-Net architecture, thereby offering better attribute control through additional learnable parameters. NeTI [30] advances this concept by proposing a neural mapper that adaptively outputs token embeddings based on the denoising timestep and specific U-Net layers. ProSpect [31] recommends optimizing multiple token embeddings tailored to different denoising timesteps based on the observation that different types of prompts, like layout, color, structure, and texture, are activated at different stages of the denoising process. Similarly, a study by [91] shows layered activation insight to learn distinct attributes by selectively activating the tokens within their respective scopes. Later, HiFiTuner [32] integrates multiple techniques into the learnable token, including mask-guided loss function, parameter regularization, time-dependent embedding, and generation refinement assisted by the nearest reference. Alternative approaches, like DreamArtist [92], choose to optimize both negative and positive prompt embeddings to refine the detail preservation. In addition to these token-level refinement approaches, the field continues to evolve with novel techniques such as InstructBooth’s reinforcement learning framework [93] and gradient-free evolutionary optimization [94]. In summary, recent developments following the foundational work of Textual Inversion focus on reducing training times and enhancing the visual quality of generated images.
@@
In the realm of TTF methods for PCS, there is a clear shift towards fine-tuning model weights rather than just the token embeddings. This approach often addresses the limitations where token embeddings alone struggle to capture complex semantics uncovered in the pre-training data [25, 95]. DreamBooth [4] proposes to use a unique modifier by a rare token to represent the SoI and fine-tune the whole parameters of the diffusion model. Besides, a regularization dataset containing 20-30 images with the same category as SoI is adopted to overcome the overfitting problem. These two combined approaches achieve impressive performance that largely promotes the progress of the research on image personalization. However, fine-tuning the entire model for each new object causes considerable storage costs, potentially rendering widespread application. To address this, Custom Diffusion [20] focuses on identifying and fine-tuning critical parameters, particularly the key-value projections in cross-attention layers, to achieve a balance of visual fidelity and storage efficiency. Further approach, Perfusion [21], also adopts the cross-attention fine-tuning and proposes to regularize the update direction of the K (key) projection towards the super-category token embedding and the V (value) projection towards the learnable token embedding. COMCAT [96] introduces a low-rank approximation of attention matrices, which drastically reduces storage requirements to 6 MB while maintaining high fidelity in the outputs. Additionally, methods like adapters [23, 24] and LoRA variants [25, 26, 27, 34, 35, 36] are increasingly utilized in personalized generation for parameter-efficient fine-tuning. It is worth noting that the token embedding fine-tuning is compatible with diffusion weight fine-tuning. Multiple methods [37, 38, 39] have started using combined weight fine-tuning.
@@
PTA framework. For practical deployment of PCS systems, fast response time is a crucial factor. PTA methods enable real-time generation (less than 10 sec/subject) by leveraging large-scale pre-training to avoid per-subject optimization during the inference phase. Re-Imagen [97] introduces a retrieval-augmented generative approach, which leverages features from text-image pairs retrieved via a specific prompt. While it is not specifically tailored for object personalization, it demonstrates the feasibility of training reference-conditioned frameworks. Later, ELITE [47] specifically targets image personalization by combining the global reference features with text embedding while incorporating local features that exclude irrelevant backgrounds. Both fused features and local features serve as conditions for the denoising process. Similarly, InstantBooth [98] retrains CLIP models to extract image features and patch features, which are injected into the diffusion model via the attention mechanism and learnable adapter, respectively. Additionally, UMM-Diffusion [61] designs a multi-modal encoder that produces fused features based on the reference image and text prompt. The text features and multi-modal hidden state are seen as guidance signals to predict a mixed noise. Another work, SuTI [52], adopts the same architecture as Re-Imagen. The difference lies in the training samples which are produced by a massive number of TTF models, each tuned on a particular subject set. This strategy promotes a more precise alignment with personalization at an instance level rather than the class level of Re-Imagen. Moreover, Domain-Agnostic [89] combines a contrastive-based regularization technique to push the pseudo embedding produced by the image encoder towards the existing nearest pre-trained token. Besides, They introduce a dual-path attention module separately conditioned on the nearest token and pseudo embedding. Compared to the methods that use separate encoders to process a single modality, some works have explored the usage of pre-trained multi-modal large language models (MLLM) that can process text and image modality within a unified framework. For example, BLIP-Diffusion [57] utilizes the pre-trained BLIP2 [62] that encodes multimodal inputs including the SoI reference and a class noun. The output embedding is then concatenated with context description and serves as a condition to generate images. Further, Customization Assistant [58] and KOSMOS-G [60] replace the text encoder of Stable Diffusion with a pre-trained MLLM to output a fused feature based on the reference and context description. Meanwhile, to meet the standard format of Stable Diffusion, a network is trained to align the dimension of the output embedding.
@@ 5.2 Personalized Style Generation
TTF framework. In this field, StyleDrop [24] leverages adapter tuning to efficiently capture the style from a single reference image. This method demonstrates the effectiveness through iterative training, utilizing synthesized images refined by feedback mechanisms, like human evaluations and CLIP scores. This approach not only enhances style learning but also ensures that the generated styles align closely with human aesthetic judgments. Later, GAL [19] proposes an uncertainty-based evaluation strategy to filter high-quality synthetic-style data and uses a weighted schema to balance the contribution of the additional samples and the original reference. Furthermore, StyleAligned [99] focuses on maintaining stylistic consistency across a batch of images. This is achieved by using the first image as a reference, which acts as an additional key and value in the self-attention layers, ensuring that all subsequent images in the batch adhere to the same stylistic guidelines. Style-friendly [100] introduces a novel diffusion model fine-tuning approach that enhances personalized artistic style generation by adaptively biasing noise sampling toward higher noise levels, where stylistic features emerge.
PTA framework. For the PTA framework, StyleAdapter [82] employs a dual-path cross-attention mechanism within the PTA framework. This model introduces a specialized embedding module designed to extract and integrate global features from multiple style references. Diptych Prompting [101] utilizes an inpainting mechanism to draw another image with the same style of the reference part.
@@ 5.3 Personalized Face Generation
Personalized face generation aims to generate diverse identity images that adhere to text prompt specifications, utilizing only a few initial face images. Compared to general object personalization, the scope is narrowed to a specific class, humans. An obvious benefit is that one can readily leverage large-scale human-centric datasets [67, 102, 103] and utilize pre-trained models in well-developed areas, like face landmark detection [104] and face recognition [105].
TTF framework. Regarding TTF methods, PromptNet [106] trains a diffusion-based network that encodes the input image and noisy latent to a word embedding. To alleviate the overfitting problem, the noises predicted by the word embedding and context description are balanced through fusion sampling in classifier-free guidance. Additionally, Celeb Basis [65] provides a novel idea that the personalized ID can be viewed as the composition of celebrity’s face, which has been learned by the pre-trained diffusion model. Based on this hypothesis, a simple MLP is optimized at test time to transform face features into the weighting of different celebrity name embeddings.
PTA framework. Thanks to the abundance of available datasets featuring the same individual in various contexts, which provide valuable data for pre-training PTA methods, the number of works in PTA frameworks is rapidly increasing. Face0 [107] crops the face region to extract refined embeddings and concatenates them with text features. During the sampling phase, the output of classifier-free guidance is replaced by a weighted combination of the noise patterns predicted by face-only embedding, text-only embedding, and concatenated face-text embedding. The $\mathcal{W}+$ Adapter [53] constructs a mapping network and residual cross-attention modules to transform the facial features from the StyleGAN [108] $\mathcal{W}+$ space into the text embedding space of Stable Diffusion. FaceStudio [109] adapts the cross-attention layer to support hybrid guidance including stylized images, facial images, and textual prompts. Moreover, PhotoMaker [46] constructs a high-quality dataset through a meticulous data collection and filtering pipeline. They use a two-layer MLP to fuse ID features and class embeddings for an overall representation of human portrait. PortraitBooth [50] also employs a simple MLP, which fuses the text condition and shallow features of a pre-trained face recognition model. To ensure expression manipulation and facial fidelity, they add another expression token and incorporate the identity preservation loss and mask-based cross-attention loss. InstantID [6] additionally introduces a variant of ControlNet that takes facial landmarks as input, providing stronger guiding signals compared to the methods that solely rely on attention fusion.
@@ 5.4 Multiple Subject Composition
TTF framework. This task presents a challenge for the TTF methods, particularly in how to integrate the parameters within the same module which are separately fine-tuned for individual SoI. Some works focus on the one-for-one generation following a fusion mechanism. For instance, Custom Diffusion [20] proposes a constrained optimization method to merge the cross-attention key-value projection weights with the goal of maximizing reconstruction performance for each subject. Mix-of-Show [25] fuses the LoRA [110] weights with the same optimization objective. StyleDrop [24] dynamically summarizes noise predictions from each personalized diffusion model. In OMG [27], the latent predicted by each LoRA-tuned model is spatially composited using the subject mask. Joint training is another strategy to cover all expected subjects. SVDiff [22] employs a data augmentation method called Cut-Mix to compose several subjects together and applies a location loss to regularize attention maps, ensuring alignment between each subject and its corresponding token. Similar strategies are found in other works [20, 65] which train a single model by reconstructing the appearance of every SoI. There are advanced control mechanisms designed to manage multiple subjects. Cones [111] proposes to find a small cluster of neurons that preserve the most information about SoI. The neurons belonging to different SoI will be simultaneously activated to generate the combination. Compositional Inversion [87] introduces spatial region assignment to different subjects to improve the composition success rate.
PTA framework. For PTA frameworks, multi-subject generation is achieved through specialized architectural designs. Fastcomposer [5], Subject-Diffusion [63], and $\lambda$-eclipse [112] place each subject feature in its corresponding placeholder in the text embedding, ensuring a seamless and efficient combination. CustomNet [113] and MIGC [114] train a PTA network that supports location control for each subject. SSR-Encoder [55] implements an encoder to selectively preserve the desired subject feature and a cross-attention module to support multi-subject feature fusion.
@@ 5.5 High-level Semantic Personalization
TTF framework. For now, all researches in this field are based on the TTF framework. ReVersion [115] intends to invert object relations from references. Specifically, they use a contrastive loss to guide the optimization of the token embedding towards specific clusters of Part-of-Speech tags, such as prepositions, nouns, and verbs. Meanwhile, they also increase the likelihood of adding noise at larger timesteps during the training process to emphasize the extraction of high-level semantic features. Lego [116] focuses on more general concepts, such as adjectives, which are frequently intertwined with the subject appearance. The concept can be learned from a contrastive loss applied to the dataset comprising clean subject images and images that embody the desired adjectives. Moreover, ADI [117] aims to learn the action-specific identifier from the references. To ensure the inversion only focuses on the desired action, ADI extracts gradient invariance from a constructed triplet sample and applies a threshold to mask out the irrelevant feature channels.
@@ 6.1 Personalization on Extra Conditions
Recent personalization tasks tend to include additional conditions for diverse content customization. One common application is to customize the subject into a fixed source image. For example, PhotoSwap [118] introduces a new task that replaces the subject in the source image with the SoI from the reference image. To address this requirement, they first fine-tunes a diffusion model on the references to obtain a personalized model. To preserve the background of the source image, they initialize the noise with DDIM inversion [119] and replace the intermediate feature maps with those derived from source image generation during inference time. Later, MagiCapture [53] broadens the scope to face customization. Another similar application can be found in Virtual Try-on, which aims to fit selected clothing onto a target person. The complexities of this task have been thoroughly analyzed in another survey [120].
Additional conditions in personalization tasks may include adjusting the layout [69], transforming sketches [121], controlling viewpoint [113, 122], or modifying poses [6]. Each of these conditions presents unique challenges and requires specialized approaches to integrate these elements seamlessly into the personalized content.